OpenStudy (anonymous):

can someone anwser this? Q&A's. √17x-60=x {5,12} 3 {20, -3} {-20, 3}

OpenStudy (kropot72):

Sorry for delay. My calculator needs a new battery. Back soon :D

OpenStudy (kropot72):

Is this the question? \[\sqrt{(17x})-60=x\]

OpenStudy (anonymous):

ya but the line is across the whole equation, if that makes a diffrence lol

OpenStudy (kropot72):

Yes. It sure makes a difference :D

OpenStudy (kropot72):

\[\sqrt{(17x-60})=x\] squaring both sides gives: x^2 - 60 = x^2 Rearranging to form a quadratic gives: x^2 - 17x + 60 = 0 This factorises to: (x - 12)(x - 5) = 0 The solutions are {5, 12} The first choice is the answer.